class 12l of 12%alcahol solutioni mixxed with a 20% alcahol solution to get a 14% solution how many liters of 20% solution are needed
Answer by stanbon(75887)

class 12L of 12% alcahol solutioni mixxed with a 20% alcahol solution to get a 14% solution how many liters of 20% solution are needed
------
alcohol + alcohol = alcohol
0.12*12L + 0.20x = 0.14(121+x)
------
Multiply thru by 100
12*12L + 20x = 14*12L + 14x
6x = 2*12L
x = (1/3)(12L)
x = 4 L (amt of 20% solution needed)
=======================================
Cheers,
Stan H.